OpenStudy (anonymous):

could someone please explain the difference between f (|x|) and |f(x)| in how it relates to graphs?

jimthompson5910 (jim_thompson5910):

f(|x|) applies the absolute value to each x value while |f(x)| applies the absolute value to the final result For instance, if f(x)=2x^2+3x+5, then f(|x|) = 2|x|^2+3|x|+5 and |f(x)| = |2x^2+3x+5| The first will be a graph that is symmetric by the y-axis while the second will always be positive (ie so the graph of |f(x)| will NOT lie in quadrants III or IV)

OpenStudy (anonymous):

so would both graphs be symmetric to the y-axis?

jimthompson5910 (jim_thompson5910):

f(|x| ) always is, but |f(x) | may not be all the time

OpenStudy (anonymous):

for \[f(|x|) \]forget what you see to the left of the y - axis . it will be the mirror image of the right. here is an example with \[f(x)=x^3-3x^2-18x+9\] http://www.wolframalpha.com/input/?i=y%3Dx^3-3x^2-18x%2B9 versus \[f(|x|)\] http://www.wolframalpha.com/input/?i=y%3D |x|^3-3|x|^2-18|x|%2B9

OpenStudy (anonymous):

and here is \[|f(x)|\] http://www.wolframalpha.com/input/?i=y%3D |x^3-3x^2-18x%2B9|

OpenStudy (anonymous):

copy and paste whole link . you will see that \[|f(x)|\] reflects what is below the y-axis to above

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3D |x^3-3x^2-18x%2B9|

OpenStudy (anonymous):

nope still didn't work. copy and paste

OpenStudy (anonymous):

i got the image of the first one but i cant get the image of |f(x)|

OpenStudy (anonymous):

for some reason the absolute values signs make it hard to copy. open wolfram and type in y=|x^3-3x^2-18x+9| you will get it

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

copying and pasting link works too, you just can't click on it

OpenStudy (anonymous):

just got it!